40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4123     Entry time: Fri Jan 7 00:49:16 2011
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Recovered Xarm Green Lock, Preparation for Beat Note Measurement 

[Kiwamu, Jenne]

We went this evening in search of a beat note signal between the Xarm transmitted green light and the PSL doubled green light. 

First, we removed our new ETMX camera (we left the mount so it should be easy to put back) from the other day.   We left the test masses exactly where they had been while flashing for IR, so even though we can no longer confirm, we expect that the IR beam axis hasn't changed.  We used the steering mirror on the end table to align the green beam to the cavity.  We turned on the loop to lock the end laser to the cavity, and achieved green lock of the arm.

Then we went to the PSL table to overlap the arm transmitted light with the PSL doubled light.  We made a few changes to the optics that take the arm transmitted light over to the PD.  We found that the arm transmitted light was very high, so we changed from having one steering mirror to having 3 (for table space / geometry reasons we needed 3, not just 2) in order to lower the beam axis.  We also found that the spot size of the arm transmitted beam was ~2x too small, so we changed the mode matching telescope from a 4x reducer to a 2x reducer by changing the 2nd lens from f=50mm to f=100mm (the first lens is f=200mm).  We made the arm transmitted beam and the PSL green beam overlap, but we saw no peak on the spectrum analyzer. 

We checked the temperature of the PSL and end lasers, and determined that we needed to adjust the set temp of the end laser.  However, we still didn't find any beat note.

We then tweaked the temperature of the doubling oven at the end station, to maximize the power transmitted, since Kiwamu said that that had worked in the past.  Alas, no success tonight.

We're stopping for the evening, with the success of reacquiring green lock of the Xarm.

ELOG V3.1.3-